OpenStudy (anonymous):

36.0kn/c=2(8.99x10^9)(Q/2.59)/.197 solve for Q

OpenStudy (jamesj):

The right hand side, RHS is \[ RHS = 2(8.99 \times 10^9)(Q/2.59)/.197 \] \[ = (3.524 \times 10^{10})Q \] \[ = \frac{36.0kn}{c} = LHS \] Now it should be easier for you.

OpenStudy (jamesj):

Yes?

OpenStudy (jamesj):

chan ... talk to me

OpenStudy (jamesj):

ok ... hope that helped. find me or someone in Physics chat if you need more help

OpenStudy (anonymous):

i just didn't know how to deal with the Q/2.59

OpenStudy (jamesj):

yes ... just put the 2.59 and the 0.197 in the denominator.

OpenStudy (anonymous):

so what did you get as a final answer.. just to compare and make sure i did it right

OpenStudy (jamesj):

I didn't calculate the final answer, but it's obviously Q = (some constant, of order -9 or so)*(kn/c)

OpenStudy (anonymous):

thanks man.. some one that finally helped me and didn't start spittin out more complicated formulas

OpenStudy (jamesj):

sure. Come back soon with some more Physics questions.